When Justin goes to work, he drives at an average speed of 55 miles per hour. It takes about 1 hour and thirty minutes for Justin to arrive at work. His car travels about 30 miles per gallon of gas. If gas costs $3.45 per gallon, how much money does Justin spend to travel each mile to work.
Answer by TimothyLamb(4379)

s = d/t
d = st
---
d = 55*1.5
d = 82.5 miles to work
---
spent per mile:
(1/30 gallon/miles) * (3.45 dollars/gallon) = 0.115
---
total spent one-way:
(82.5 miles) * (1/30 gallon/miles) * (3.45 dollars/gallon) = 9.4875
---
answer:
expense per mile = $0.115
expense one-way to work = $9.49
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php